What is the acceleration of a meteor at a distance 8*10^6m from the center of the earth?
1
Expert's answer
2019-07-15T09:10:04-0400

Using Newton's law of gravitation


"F_{gravitational}=G*M_{earth}*m_{meteor}\/r^2"

Using Newton's law


"F=m*a"

"a_{meteor}=(G*M_{earth}*m_{meteor}\/r^2)\/m_{meteor}="

"a_{meteor}=(G*M_{earth}\/r^2)="

"6.674*10^{-11}*5.972*10^{24}\/(8*10^6)^2=6.22768 m\/s^2"

Answer: 6.22768 m/s2.